OpenStudy (roadjester):

\[\huge{\int\limits\limits\limits\limits_{0}^{{\frac \pi 2}}}\int\limits\limits\limits\limits_{0}^{\cos \theta} e^{\sin \theta}d r d \theta\]

OpenStudy (anonymous):

If you integrate first with respect to r, you will be able to use a nice substitution afterwards. \[ u= sin \theta\]

OpenStudy (unklerhaukus):

\[{\int\limits_{0}^{{\frac \pi 2}}}\int\limits\limits\limits\limits_{0}^{\cos \theta} e^{\sin \theta}d r d \theta\]\[={\int\limits_{0}^{{\frac \pi 2}}}e^{\sin \theta}\int\limits_{0}^{\cos \theta} d r d \theta\]

OpenStudy (roadjester):

uh, can you put "huge" in front?

OpenStudy (roadjester):

it makes the text larger

OpenStudy (unklerhaukus):

\[=\huge{\int\limits_{0}^{{\frac \pi 2}}}e^{\sin \theta}\int\limits_{0}^{\cos \theta} d r d \theta\]

OpenStudy (unklerhaukus):

\[=\huge{\int\limits_{0}^{{\frac \pi 2}}}e^{\sin \theta}\left(r|_{0}^{\cos \theta}\right) d \theta\]

OpenStudy (anonymous):

I get A=e-1

OpenStudy (unklerhaukus):

\[=\huge{\int\limits_{0}^{{\frac \pi 2}}}e^{\sin \theta}{\cos \theta} d \theta\]

OpenStudy (anonymous):

Um, this equation looks like it's missing an r. usually double integrals of this type are integrated with respect to r*dr*d(theta)

OpenStudy (unklerhaukus):

that is a good point @agentc0re

OpenStudy (roadjester):

The "r" wasn't in the question. It's question 5.

OpenStudy (anonymous):

Guess it's been left away so the substitution becomes more simple.

OpenStudy (roadjester):

And I don't even know what was done...

OpenStudy (roadjester):

I'm starting to despise double integrals and I just started.

OpenStudy (anonymous):

lol. just wait until you do triple integrals and have to change your bounds of integration. :D Honestly it's not too bad. My guess is that it's not a true volume oriented problem so there is no r*dr*d(theta). Looks like you were doing good before though, keep it up!

OpenStudy (unklerhaukus):

\[{\int\limits_{0}^{{\frac \pi 2}}}e^{\sin \theta}{\cos \theta} d \theta\] \[\text{let } u=\sin \theta;\qquad\text du=\cos\theta \text d\theta\] \[\theta =\frac{\pi}{2}\longrightarrow u=1\]\[\theta =0\longrightarrow u=0\] \[={\int\limits_{0}^{1}}e^{u}\text du\]

OpenStudy (roadjester):

@agentc0re uh huh, right. Keep up with the sarcasm and I'll find an Urk-Khai!

OpenStudy (unklerhaukus):

can you finish from here/?

OpenStudy (roadjester):

So the answer would be e-1?

OpenStudy (anonymous):

yes @roadjester